OpenStudy (anonymous):

x2 – x = 2 a) x = -1, x = -2 b) x = -2, x = 1 c) x = 2, x = -1 d) x = 1, x = 2

OpenStudy (anonymous):

x(x-1)=2

OpenStudy (anonymous):

c) x = 2, x = -1

myininaya (myininaya):

x^2-x-2=0 (x-2)(x+1)=0 x=2 and x=-1 are the solution

OpenStudy (anonymous):

thanks!